) if the supply curve is q=3+1.5p what is the producer surplus if p=12?

1
Expert's answer
2021-11-03T15:49:31-0400

Supply curve: "q=3+1.5p\\p=\frac{q-3}{1.5}=0.67q-2"

"-q( at\space p=12)=3+1.5(12)=21\space units"

q at (p=0) = 3 units




"P.S=\frac{1}{2}(market\space price-min\space price\space willing\space to\space receive)\times q\\=\{(12\times 3)\}+\{\frac{1}{2}(21-3)\times12\}\\=\$144"